OpenStudy (anonymous):

Evaluate the integrals in simplest way possible ∬P (dot)n ds over upper half of the sphere r = 1 if P = curl (jx-kz) but if divergence theorem is used del cross del dot y = 0 y=(jx-kz) so whats the possible way to solve this without using divergence theorem? pls help...

OpenStudy (anonymous):

pls help me with thgis question..... i dont get this, how will i express the volume integral in double integrals......

OpenStudy (anonymous):

Curl (0,x,-z) = (0,0,1) n ds = (x/z, y/z, 1) dA ∬ dA where A is the disk x^2 + y^2 <= 1 = pi but its a sphere, how can i write x^2 + y^2 <= 1